# Cluster Haptic Texture Database: Haptic Texture Database with Varied Velocity–Direction Sliding Contacts

### v2

@@ Methods
Experiment II: Texture Classification with Motion Labels
We examined how the velocity and direction of sliding contact affect the classification accuracy of texture identification. Previous studies have shown that human perception of texture is significantly influenced by the motion parameters during tactile exploration [2, 3]. Building upon these findings, we investigate whether incorporating motion information (velocity and direction) as additional features can enhance the performance of our texture classification system, similar to how humans integrate motion cues in their texture perception.
Experiment Details
We used the CNN (ResNet-34) model that showed the best overall performance in Experiment I as our baseline model, and performed texture classification using the multimodal architecture shown in Fig. 5. We trained the classifier using audio signals, acceleration signals, and combined audio-acceleration signals, with and without motion labels (velocity and direction) as additional features. We then compared the classification accuracies between these conditions to investigate how motion information affects texture classification performance.
Results and Discussion
The improvement in classification accuracy when adding motion labels to acceleration signals suggests that explicit motion information helps the classifier focus on texture-specific features in the acceleration data. This phenomenon might be analogous to how humans integrate motion cues in texture perception, where motion parameters can influence the perception of surface properties. However, the limited improvement in accuracy for audio-based and combined audio-acceleration classifiers indicates that audio signals already contain sufficient texture-specific information, making additional motion cues less beneficial. This could be due to the inherent stability of audio signals across different motion parameters, which may not significantly alter the texture-specific characteristics captured by the audio modality. Furthermore, the high sampling rate of audio data might capture features beyond human perceptual capabilities, suggesting that the relationship between motion and texture perception in our system may differ from human perception.
Experiment III: Velocity Classification
We classified five velocities to evaluate how variations in the velocity of tracing affect the sensor signal in our haptic texture database.
Experiment Details
We used a CNN-based classifier for the velocity classification task, as it achieved the highest classification accuracy in Experiment I. Unlike in Experiment I, where we used texture material IDs as labels, we used five different traversal velocities as labels in this experiment. We trained a separate classifier for each texture material and calculated the classification accuracy of the velocity labels for each one to consider the results in terms of the physical properties of the textures.
Results and Discussion
The classification accuracies of velocity labels by texture for classifiers using audio signals, acceleration signals, and both audio and acceleration signals are shown in Fig. 8a-c, respectively. The accuracies are highest when both audio and acceleration signals are used, similar to Experiment I, with an average accuracy of 88.76 %. Interestingly, unlike the texture classification results in Experiment I, classifiers using only acceleration signals outperformed those using only audio signals in the velocity classification task.
On the other hand, materials with low accuracies, such as cotton gauze and goat skin (Fig. 9b), possess challenging physical properties for velocity classification. Cotton gauze is thin and soft with a high coefficient of friction against the artificial fingertip, while goat skin features a notably rough and irregular surface. These materials’ non-uniform structures create inconsistent mechanical interactions during traversal at different velocities. Consequently, the sensor signals exhibit unpredictable patterns that do not correlate systematically with velocity changes, making it difficult for the classifier to identify distinctive velocity-dependent features in the data.
Nevertheless, this dataset was collected with an artificial fingertip, and the above discussions may not translate perfectly to interactions involving a real human finger. Even with this limitation, the combined audio–acceleration model still achieves a mean classification accuracy of 88.76 %, substantially exceeding chance level, which indicates that the key signal features are captured robustly by the artificial fingertip. Consequently, supplementing this dataset with recordings obtained from actual human fingertips [27, 25] should facilitate analyses and learning schemes that more closely mirror human tactile perception.
Experiment IV: Direction Classification
To clarify how changes in touch direction affect the signal captured by the sensor in our tactile texture database, we performed a classification task to classify eight different touch directions.
@@
Experiment Details
For the direction classification, we used the same CNN-based classifier as in Experiment II. In this experiment, eight different traversing directions were set as labels to be classified, and a separate classifier was trained on each texture separately to compute the classification accuracy of the direction labels.
Results and Discussion
The classification accuracies of direction labels by texture for classifiers using audio signals, acceleration signals, and both audio and acceleration signals are shown in Fig. 10a-c, respectively. Compared to the velocity label classification in Experiment II, the accuracy tended to be lower for all classifiers using any data in the direction label classification. The classifier with the highest average accuracy was the one that used both audio and acceleration signals, with an average accuracy of 78.79 %. The classification accuracy using the acceleration signal was higher than the audio-based model, similar to the results of Experiment II.
The materials were also examined separately to discuss the physical properties of the results in terms of direction classification accuracy. A list of texture material types with high and low accuracy and their accuracy for classifiers using both audio and acceleration signals is shown in Table 6. From Table 6, it can be seen that the textures with the highest accuracy were gumwood and north pipe (Fig. 11a). These woods exhibit pronounced, quasi-periodic wood grain structures along a dominant orientation. When the artificial finger traverses the surface, the number of wood grain ridges crossed per unit time changes systematically with scan direction, producing direction-specific stick-slip events reflected in the acceleration and acoustic spectra.

### v4

@@ Methods
We classified five velocities to evaluate how variations in the velocity of tracing affect the sensor signal in our haptic texture dataset.
We used a CNN-based classifier for velocity classification, as it showed the best performance in Experiment I. In this experiment, five velocities were used as labels instead of texture IDs. For each texture, we trained a separate classifier and calculated the velocity classification accuracy to examine the influence of texture properties.
Fig. 9a-c shows the velocity label classification accuracies by texture for each classifier. The accuracies are highest when both audio and acceleration signals are used, similar to Experiment I, with an average accuracy of 88.76 %.
On the other hand, materials with low accuracies, such as cotton gauze and goat skin (Fig. 10b), possess challenging physical properties for velocity classification. Cotton gauze is thin and soft with a high coefficient of friction against the rubber tip, while goat skin features a notably rough and irregular surface. These materials’ non-uniform structures create inconsistent mechanical interactions during traversal at different velocities. Consequently, the sensor signals exhibit unpredictable patterns that do not correlate systematically with velocity changes, making it difficult for the classifier to identify distinctive velocity-dependent features in the data.
Experiment IV: Direction Classification
@@
To clarify how changes in sliding direction affect the signal captured by the sensor in our tactile texture dataset, we performed a classification task to classify eight different sliding directions.
For direction classification, we used the same CNN-based classifier as in Experiment III. Eight traversing directions were set as labels, and classification accuracy was computed for each texture.
Fig. 11a-c shows the direction label classification accuracies by texture for each classifier. Compared to the velocity label classification in Experiment III, the accuracy tended to be lower for all classifiers using any data in the direction label classification. The classifier with the highest average accuracy was the one that used both audio and acceleration signals, with an average accuracy of 78.79 %.
A list of texture material types with high and low accuracy and their accuracy for classifiers using both audio and acceleration signals is shown in Table 7. Notably, the textures with the highest accuracy were gumwood and north pipe (Fig. 12a). These woods exhibit pronounced, quasi-periodic wood grain structures along a dominant orientation. When the rubber tip traverses the surface, the number of wood grain ridges crossed per unit time changes systematically with scan direction, producing direction-specific stick-slip events reflected in the acceleration and acoustic spectra.